OpenStudy (anonymous):

i neep help babdly.. going back to college after years of being out.. have a placement test today.. I need help with Polynomials

OpenStudy (anonymous):

\[(x+y)^{2} - (9xy-6x ^{2})=\]

OpenStudy (anonymous):

someone please walk me thru this

OpenStudy (ash2326):

@deemagz12 do you know what's \((a+b)^2\)??

OpenStudy (anonymous):

\[a ^{2} +b ^{2}\]

OpenStudy (fwizbang):

I know you're anxious, but the placement test is just a way of getting you into classes that are the right level so you can be successful. Just relax and do your best ....

OpenStudy (anonymous):

i am aware mi still want to do well

OpenStudy (fwizbang):

About the worst outcome of a placement test is for you to be placed in classes that you're not ready for. I'm still happy to help you though. Do you know FOIL?

OpenStudy (anonymous):

first outer inner last.. but I didnt think I would have to use FOIL for this

OpenStudy (fwizbang):

Use it to calculate (a+b)^2=(a+b)*(a+b)

OpenStudy (anonymous):

no idea how this